### v1

@@ IX-B Optimization problem of the central controller with autonomous incentivized LLMAs
$\displaystyle\mathcal{H}_{k}=\sigma(\{\pi_{0},u_{0},\dots,u_{k},p_{1},\dots,p_{k-1}\})$
With the estimates of the cost of the individual LLMAs and observation matrices, the central controller can use $\mathcal{H}_{k}$ to compute the public belief using (15). If the cost and observation matrices are not available, the central control can learn the optimal policy using stochastic approximation, as discussed in the next section. Similar to the previous subsection, $\pi_{k}$ can be shown to be a sufficient statistic for the filtration $\mathcal{H}_{k}$ and therefore the incentivization policy $\mu:\mathcal{P}(\mathcal{X})\to[0,1]$ of the central controller determines the incentives as,
$\displaystyle p_{k+1}=\nu(\pi_{k})$
@@
Although classical methods like value iteration can be used to solve the continuous-valued optimization problem of (26), for the simple case of $|\mathcal{U}|=|\mathcal{X}|=2$, we show in the next subsection that the optimal policy has a threshold structure which can be searched much more efficiently using stochastic approximation. Note that this reduction is still practical for failure state detection e.g. bad product identification using product reviews on online platforms.
@@ IX-C Structural Results
$\displaystyle\chi(y,\pi)=\frac{\alpha_{2}-\alpha_{1}}{\omega_{2}-\omega_{1}}\frac{B_{y}\pi}{\mathds{1}B_{y}\pi}+\frac{\Delta_{2}-\Delta_{1}}{\omega_{2}-\omega_{1}}$ (27)
@@
Under (S4) and (S5), the optimal incentive policy of (26) is a threshold with respect to the public prior and can be computed as,
$\displaystyle\nu^{*}(\mu)=\begin{cases}0,&\text{if}\ \pi(2)\in[0,\bar{\pi}]\\ \chi(y,\pi),&\text{if}\ \pi(2)\in[\bar{\pi},1],\end{cases}$ (28)
@@
Similar to the result of Theorem 3, this theorem shows that the optimal incentivization policy of the central controller is threshold in the public prior. Since the state space has cardinality 22 22, the threshold switching curve becomes a single threshold point $\bar{\pi}$, and can be efficiently searched.
@@
Further, the incentivization function of (27) and an optimal policy with a threshold structure of (28) implies that the incentive sequence by the central controller is a sub-martingale. This result and a concentration inequality type bound on the total incentive spent are formalized in the next result.
@@
Considered the controlled incentivized fusion of information from LLMAs where the cost function is (25) and the optimal incentive policy, $\nu^{*}$ satisfies (26) with the incentive function of (27). Under (A1) the optimal incentive sequence $p_{k}=\nu^{*}(\pi_{k-1})$ is a sub-martigale, i.e., $p_{k}\geq\mathds{E}\{p_{k-1}\}$. Further, the total incentive spent is such that in a sample path is such that, $\mathds{P}(\sum_{1\leq k\leq T}p_{k}\geq B)\leq\frac{T}{B}$, where $B$ can be considered as a budget constraint.
The above theorem characterizes the nature of a sample path of incentivization and secondly provides a bound on the probability that the total incentive exceeds the budget $B$. This helps analyze the deviation of the total expenditure from the budget $B$ which is a constraint of the central controller.
@@ IX-D Summary
Autonomous LLMAs are already provided by third-party services that offer them as a service [66]. They have a unit monetary cost associated with using them for any application. Motivated by such LLMAs, we studied stochastic control of autonomous LLMAs who are incentivized by a central controller to perform Bayesian inference. We showed structural results on the optimal incentive policy and a concentration inequality, which characterized the probability that the central controller would exceed their budget. The next section proposes a stochastic approximation algorithm that exploits the structural results of Theorem 3 and Theorem 4 to search for the corresponding threshold policy of the socialistic LLMA and the central controller, respectively.
X Stochastic approximation for estimating the optimal policy to control herding in LLMAs
We propose a stochastic approximation algorithm that searches for the optimal policy for the stochastic control of the socialistic LLMAs (21) and the central controller (26), which have the threshold structure of (22) and (28), respectively.
@@ X Stochastic approximation for estimating the optimal policy to control herding in LLMAs
When LLMAs are deployed on online platforms to perform Bayesian inference for various purposes, one needs to control for the herding the LLMAs exhibit. For this purpose, we proposed two optimal stopping formulations for centrally controlled and autonomous LLMAs. The efficiency of estimation of the optimal policy parameters is especially important in real-life applications where the time to update the policy parameters is limited. Another constraint is that in real-life systems, access to the observation matrices is limited; therefore, the estimation has to be done without the knowledge of the system parameters. Therefore, we use structural results on the optimal policies for the stopping time formulations and use stochastic approximation to estimate the optimal policy.
X-B Stochastic Approximation Algorithm
Searching for a hard threshold of the form  (22) and (28) using a discrete search is a computationally intesive task. We, therefore, propose approximating the hard threshold policy by a sigmoidal policy of the form,
@@ X-B Stochastic Approximation Algorithm
where $\Theta$ is the policy parameter representing the threshold, and $\varepsilon$ is the approximation parameter, and the policy converges to a hard threshold policy at $\Theta$ as $\varepsilon\to 0$. To restrict the policy parameter to be in the range $[0,1]$, we can reparameterize it as $\sin^{2}(\theta)$. The approximate parameterized policy of (29) can be used to obtain an approximate value of the cumulative value function (of either (25) or (19)) by interacting with the system using the policy for $T_{e}$ interactions.
For the socialistic cost of (19), we can compute the approximate cost with respect to a policy by
$\displaystyle\begin{split}&\hat{J}(\hat{\mu}(\cdot,\Theta_{k}))\\ =&\sum_{k=1}^{T_{e}}\hat{\mu}(\pi_{k},\Theta_{k})\left(\rho^{k-1}c(x,u_{k})+\rho^{k-1}d\mathds{1}(x=e_{1})\right)+\\ &\ \rho^{T_{e}-\sum_{k=1}^{T_{e}}\lfloor 0.5+\hat{\mu}(\pi_{k},\Theta_{k})\rfloor}\left[\Upsilon\mathds{1}(x\neq e_{1})+\ \frac{\min_{u\in\mathcal{U}}c(x,u)}{1-\rho}\right],\end{split}$ (30)
where $\sum_{k=1}^{T_{e}}\lfloor 0.5+\hat{\mu}(\pi_{k},\Theta_{k})\rfloor$ just computes the empirical stopping time in place of the stopping time $\tau$ in (19). Note that in the above equation, the filtration and the expectations have been replaced with the realized cost, and hence, this is a noisy estimate of the true expected cost of (31).
@@
$\displaystyle\hat{J}(\hat{\mu}(\Theta_{k}))=\sum_{k=1}^{T_{e}}f(p_{k},k).$ (31)
We now describe our simultaneous pertubation based stochastic approximation algorithm which is summarized in Algorithm 7. We perform $H$ learning episodes, each of which is initialized with the same conditions, and we adaptively update the policy parameter $\Theta_{m}$ across these episodes. For each episode $m$, we perturb the policy parameter by ±$\pm\delta$ and obtain two policy parameters $\Theta^{+}_{m}$ and $\Theta^{-}_{m}$. We then approximate the cost function for both the parameters using (31) by performing $T_{e}$ interactions each. We approximate the gradient by finite-difference method (step 6 of Algorithm 7). Finally, the policy parameter uses a gradient descent step with a step size $\Theta_{m}$. If the parameters of the system are known to be more or less constant, the step size is decreasing to ensure asymptotic convergence [44], else a constant step-size can be used to track changes in the true threshold parameter.
1:Input: Initial Parameter $\Theta_{0}$, Perturbation $\delta$, Time Horizon $H$, Step Sizes $(\beta_{m})$, Episode Length $T_{e}$
@@
5:     Approximate cost with (31) using $T_{e}$ interactions with $\Theta_{k}^{+}$ and $\Theta_{k}^{-}$, $\hat{J}(\hat{\mu}(\Theta_{k}^{+}))$ and $\hat{J}(\hat{\mu}(\Theta_{k}^{-}))$.
@@
Algorithm 7 Stochastic approximation for estimating optimal policy
@@
The above algorithm does not need access to the probability distributions $B$, which is expensive to obtain, especially in the incentivized case. Also, if the underlying parameters of the setup are evolving on a slower timescale, the above algorithm can be run with a constant step size to make the policy parameters track the changes in the system. These are the key advantages of using Algorithm 7 compared to the value iteration or other sub-optimal methods for solving our proposed non-standard POMDPs. The computational complexity of Algorithm 7 for each iteration is $\mathsf{O}(T_{e})$.
@@ X-C Summary
This section proposed a computationally efficient stochastic approximation algorithm to estimate the optimal policy of the stopping time problems for LLMAs of Section VIII and Section IX. This algorithm does not need the underlying parameters of LLMAs and is adaptive to changes to the underlying parameters. The next section presents numerical studies demonstrating the efficacy of our proposed model.
@@ XI-A Motivation
We first describe the two main real-world tasks and datasets on which the numerical results are presented. Then, we show how our construction of a single LLMA leads to interpretability. We extend the exemplary study from Section IV and conduct more extensive experiments. Next, we show numerical results on herding phenomena in a sequence of LLMAs. Finally, we show how the optimal policy for optimal stopping has a threshold structure. We also show the efficacy of stochastic approximation in a simulated setting.
@@ XI-B Task Description
For both tasks, the LLMA is an online detection mechanism equipped with a large language model (LLM) sensor that analyses comments and flags the state sequentially. The LLM is used to parse the text and obtain a list of appropriate features from the finite-dimensional feature space $\mathcal{Y}$, which the platform could design. The features contain information about the text and comparisons of the text with a given context. In our setup, the Bayesian engine of the LLMA consists of a likelihood parameterized by a neural network. For a discrete distribution, the likelihood neural network could use a restricted Boltzmann machine (RBM) to generate samples from the likelihood with 10001000 10001000 samples. The experiment results are averaged over $N_{MC}$ independent runs mentioned along with each experiment. The posterior can be updated using (13) using the likelihood and the closed form prior of (15) from the previous step. The LLMA takes an action according to (14). The action is classifying whether the user has hateful intent, and the cost accounts for the misclassification of the state. We use Mixtral 7B, an open-source mixture of experts LLM with 7 billion parameters [33], Llama-3 [74] with 70b parameters, and ChatGPT-4o which is a closed source LLM. The details of which LLM was used for this experiment is in the Appendix. We query using the TogetherAI API for the open-source LLMs (Mixtral and Llama) and OpenAI API for ChatGPT.
@@ XI-B1 Hate Speech Classification
Motivation: Flagging users who spread toxic content online is a significant challenge. The state $x_{k}$ could represent the category of peddlers classified based on the intensity and type of content they are propagating. For example, the state could be 3-dimensional, indicating the hate-intent of the user (hateful or not), the hate speech intensity scale [5], and the particular group the hate speech is directed towards. The noisy observations are the text comments from the user that inform about the state and are from a high-dimensional observation space.
In a social network, there may be multiple LLMA deployed to flag malicious users and decrease the propagation of hate speech. The flags by the previous LLMA are visible, but the private observations are not due to computation and privacy restrictions (so that the LLM can not be fine-trained on the text or the feature mappings). Since the observations are generated sequentially, the LLMA may use the same LLM but with a different context and for a different text observation. Therefore, a single LLMA can be viewed as a sequence of LLMAs learning from their private observations and the past actions of previous LLMAs.
@@
The first time accounts for the misclassification of a hateful user, and the second term accounts for the difference in intensity. We use the measuring hate speech dataset from [65], which contains 4000040000 4000040000 annotated comments from Twitter, Reddit, and Gab. The annotations are performed by crowdsourcing and indicate if the comments contain hate speech and the intensity of the toxicity exhibited on a scale of 1 to 5, measured using a Rasch measurement scale [65].
@@ XI-B2 Product Quality Identification
In the case of an e-commerce platform, the restriction on the sharing of the private observation is motivated by the compute constraints that the LLMAs will have when performing Bayesian inference on millions of products and billions of reviews. The public prior can be efficiently computed based on the output (action) of each of the LLMA.
@@
We now detail on what we mean by quality and the definition of the states. Since we don’t have access to any additional dataset other than ratings, we consider products with at least 10001000 10001000 ratings and compute the quality state ex-ante using these ratings. Therefore, the quality here is more representative of the perceived value for money [82]. We consider the following ranges $[0,3.3),[3.3,5]$ of the average review of all the ratings to assign a product quality as bad, medium, and good, respectively. To clarify, we just use all 10001000 10001000 ratings to compute the state; the observations are sampled one at a time and analyzed by the LLMA.
@@
Figure 12: Interpretability of the outputs of an LLM Sensor for Product Quality Identification. There are 4 different features that the LLM of the LLMA extracts. It can be seen that the features, indeed, are consistent with the overall rating provided by the reviewer for the two different kinds of products and can be used for a more fine-grained interpretable analysis.

### v2

@@ IX-B Optimization problem of the central controller with incentivized LLMAs
$\displaystyle\mathcal{H}_{k}=\sigma(\{\pi_{0},u_{0},\dots,u_{k},p_{1},\dots,p_{k-1}\}).$
With the estimates of the cost of the individual LLMAs and observation matrices, the central controller can use $\mathcal{H}_{k}$ to compute the public belief using (15). If the cost and observation matrices are not available, the central control can learn the optimal policy using policy gradient algorithm, as discussed in the next section. Similar to the previous subsection, $\pi_{k}$ can be shown to be a sufficient statistic for the filtration $\mathcal{H}_{k}$ and therefore the incentivization policy $\nu:\mathcal{P}(\mathcal{X})\to\mathbb{R}^{+}\cup\{0\}$ of the central controller determines the incentives as,
$\displaystyle p_{k+1}=\nu(\pi_{k}).$
@@
Although classical methods like value iteration can be used to solve the continuous-valued optimization problem of (26), for the simple case of $|\mathcal{U}|=|\mathcal{X}|=2$, we show in the next subsection that the optimal policy has a threshold structure which can be searched much more efficiently using a policy gradient algorithm. Note that this reduction is still practical for failure state detection e.g. bad product identification using product reviews on online platforms.
@@ IX-C Structural Results
$\displaystyle\chi(y,\pi)=\frac{\alpha_{2}-\alpha_{1}}{\omega_{2}-\omega_{1}}\frac{B_{y}\pi}{\mathds{1}B_{y}\pi}+\frac{\Delta_{2}-\Delta_{1}}{\omega_{2}-\omega_{1}}.$ (27)
@@
Under (S4) and (S5), the optimal incentive policy $\nu^{*}:\mathcal{P}(\mathcal{X})\to\mathbb{R}^{+}\cup\{0\}$ of (26) is a threshold with respect to the public belief and can be computed as,
$\displaystyle\nu^{*}(\mu)=\begin{cases}0,&\text{if}\ \pi(2)\in[0,\bar{\pi}]\\ \chi(y,\pi),&\text{if}\ \pi(2)\in[\bar{\pi},1]\end{cases},$ (28)
@@
Similar to the result of Theorem 3, this theorem shows that the optimal incentivization policy of the central controller is threshold in the public belief. Since the state space has cardinality 22 22, the threshold switching curve becomes a single threshold point $\bar{\pi}$, and can be efficiently searched.
@@
Further, the incentivization function of (27) and an optimal policy with a threshold structure of (28) implies that the incentive sequence by the central controller is a sub-martingale. This result and a concentration inequality type bound on the cumulative incentive spent are formalized in the next result.
@@
Considered the controlled incentivized fusion of information from LLMAs where the cost function is (25) and the optimal incentive policy, $\nu^{*}$ satisfies (26) with the incentive function of (27). Under (A1) the optimal incentive sequence $p_{k}=\nu^{*}(\pi_{k-1})$ is a sub-martingale, i.e., $p_{k}\geq\mathds{E}\{p_{k-1}\}$. Further, the cumulative incentive spent is such that in a sample path is such that, $\mathds{P}(\sum_{1\leq k\leq T}p_{k}\geq B)\leq\frac{T}{B}$, where $B$ can be considered as a budget constraint.
The above theorem characterizes the nature of a sample path of incentivization and secondly provides a bound on the probability that the cumulative incentive exceeds the budget $B$. This helps analyze the deviation of the total expenditure from the budget $B$, which is a constraint of the central controller.
@@ IX-D Summary
Autonomous LLMAs are already offered by third-party entities as a service [17, 87, 7]. They have a unit monetary cost associated with using them for any application. Motivated by such LLMAs, we studied stochastic control of autonomous LLMAs who are incentivized by a central controller to perform Bayesian inference. We showed structural results on the optimal incentive policy and a concentration inequality, which characterized the probability that the central controller would exceed their budget. The next section proposes a policy gradient algorithm that exploits the structural results of Theorem 3 and Theorem 4 to search for the corresponding threshold policy of the social welfare LLMA and the central controller, respectively.
X Policy Gradient for estimating the optimal policy to control herding in LLMAs
We propose a policy gradient algorithm that searches for the optimal policy for the stochastic control of the social welfare LLMAs (21) and the central controller (26), which have the threshold structure of (22) and (28), respectively.
@@ X Policy Gradient for estimating the optimal policy to control herding in LLMAs
When LLMAs are deployed on online platforms to perform Bayesian inference for various purposes, one needs to control for the herding the LLMAs exhibit. For this purpose, we proposed two optimal stopping formulations for centrally controlled and autonomous LLMAs. The efficiency of estimation of the optimal policy parameters is especially important in real-life applications where the time to update the policy parameters is limited. Another constraint is that in real-life systems, access to the observation matrices is limited; therefore, the estimation has to be done without the knowledge of the system parameters. Therefore, we use structural results on the optimal policies for the stopping time formulations and use policy gradients to estimate the optimal policy.
X-B Policy Gradient Algorithm
Searching for a hard threshold of the form  (22) and (28) can be formulated as a discrete stochastic optimization problem. However, in this section, we relax the problem to a continuous stochastic optimization by approximating the hard threshold policy by a sigmoidal of the form,
@@ X-B Policy Gradient Algorithm
where $\Theta$ is the policy parameter representing the threshold, and $\varepsilon\in(0,1]$ is the approximation parameter, and the policy converges to a hard threshold policy at $\Theta$ as $\varepsilon\to 0$. To restrict the policy parameter to be in the range $[0,1]$, we can reparameterize it as $\sin^{2}(\theta)$. The approximate parameterized policy of (29) can be used to obtain an approximate value of the cumulative value function (of either (25) or (19)) by interacting with the system using the policy for $T$ interactions.
For the social welfare cost of (19), we can compute the approximate cost with respect to a policy by
$\displaystyle\begin{split}&\hat{J}(\hat{\mu}(\cdot,\Theta_{k}))\\ =&\sum_{k=1}^{T}\hat{\mu}(\pi_{k},\Theta_{k})\left(\rho^{k-1}c(x,u_{k})+\rho^{k-1}d\mathds{1}(x=e_{1})\right)+\\ &\ \rho^{T-\sum_{k=1}^{T}\lfloor 0.5+\hat{\mu}(\pi_{k},\Theta_{k})\rfloor}\left[\Upsilon\mathds{1}(x\neq e_{1})+\ \frac{\min_{u\in\mathcal{U}}c(x,u)}{1-\rho}\right],\end{split}$ (30)
where $\sum_{k=1}^{T}\lfloor 0.5+\hat{\mu}(\pi_{k},\Theta_{k})\rfloor$ just computes the empirical stopping time in place of the stopping time $\tau$ in (19). Note that in the above equation, the filtration and the expectations have been replaced with the realized cost, and hence, this is a noisy estimate of the true expected cost of (31).
@@
$\displaystyle\hat{J}(\hat{\mu}(\Theta_{k}))=\sum_{k=1}^{T}f(p_{k},k).$ (31)
We now describe our simultaneous pertubation based policy gradient algorithm, which is summarized in Algorithm 7. We perform $H$ learning episodes, each of which is initialized with the same conditions, and we adaptively update the policy parameter $\Theta_{m}$ across these episodes. For each episode $m$, we perturb the policy parameter by ±$\pm\delta$ and obtain two policy parameters $\Theta^{+}_{m}$ and $\Theta^{-}_{m}$. We then approximate the cost function for both the parameters using (31) by performing $T$ interactions each. We approximate the gradient by finite-difference method (step 6 of Algorithm 7). Finally, the policy parameter uses a gradient descent step with a step size $\Theta_{m}$. If the parameters of the system are known to be more or less constant, the step size is decreasing to ensure asymptotic convergence [88], else a constant step-size can be used to track changes in the true threshold parameter.
1:Input: Initial Parameter $\Theta_{0}$, Perturbation $\delta$, Time Horizon $H$, Step Sizes $(\beta_{m})$, Episode Length $T$
@@
5:     Approximate cost with (31) using $T$ interactions with $\Theta_{k}^{+}$ and $\Theta_{k}^{-}$, $\hat{J}(\hat{\mu}(\Theta_{k}^{+}))$ and $\hat{J}(\hat{\mu}(\Theta_{k}^{-}))$.
@@
Algorithm 7 Stochastic gradient algorithm for estimating optimal policy
@@
The above algorithm does not need access to the probability distributions $B$, which is expensive to obtain, especially in the incentivized case. Also, if the underlying parameters of the setup are evolving on a slower timescale, the above algorithm can be run with a constant step size to make the policy parameters track the changes in the system. These are the key advantages of using Algorithm 7 compared to the value iteration or other sub-optimal methods for solving our proposed non-standard POMDPs. The computational complexity of Algorithm 7 for each iteration is $\mathsf{O}(T)$.
Remark 23.
For the centrally controlled case, only a single parameter needs to be communicated across the LLMAs to learn the optimal stopping time policy. Further, since the LLMAs act in a sequential fashion, such a parameter can be communicated either by the central controller or by the previous LLMA, making this a scalable approach. For the incentivization policy too, a single parameter needs to be estimated by the central entity employing the autonomous LLMAs, which makes the policy estimation protocol independent of the number of agents, and therefore scalable.
@@ X-C Summary
This section proposed a computationally efficient policy gradient algorithm to estimate the optimal policy of the stopping time problems for LLMAs of Section VIII and Section IX. This algorithm does not need the underlying parameters of LLMAs and is adaptive to changes to the underlying parameters. The next section presents numerical studies that illustrate how these models can be used to control LLMAs.
@@ XI-A Motivation
We first describe the two main real-world tasks and datasets on which the numerical results are presented. Then, we show how our construction of a single LLMA leads to interpretability. We extend the exemplary study from Section IV and conduct more extensive experiments. Next, we show numerical results on herding phenomena in a sequence of LLMAs. Finally, we show how the optimal policy for optimal stopping has a threshold structure. We also show the efficacy of policy gradient algorithm in a simulated setting.
@@ XI-B Task Description
For both tasks, the LLMA is an online detection mechanism equipped with a large language model (LLM) sensor that analyses comments and flags the state sequentially. The LLM is used to parse the text and obtain a list of appropriate features from the finite-dimensional feature space $\mathcal{Y}$, which the platform could design. The features contain information about the text and comparisons of the text with a given context. In our setup, the Bayesian engine of the LLMA consists of a likelihood parameterized by a neural network. For a discrete distribution, the likelihood neural network uses restricted Boltzmann machine (RBM) to generate samples from the likelihood with 10001000 10001000 samples of annotated comments (paragraphs). The experiment results are averaged over $N_{MC}$ independent runs mentioned along with each experiment. The posterior can be updated using (13) using the likelihood and the closed form prior of (15) from the previous step. The LLMA takes an action according to (14). The action is classifying whether the user has hateful intent, and the cost accounts for the misclassification of the state. We use Mixtral 7B, an open-source mixture of experts LLM with 7 billion parameters [57], LLaMA-3 [58] with 70b parameters, and ChatGPT-4o which is a closed source LLM. The details of which LLM was used for this experiment is in the Appendix. We query using the TogetherAI API for the open-source LLMs (Mixtral and LLaMA) and OpenAI API for ChatGPT.
@@ XI-B1 Hate Speech Classification
Motivation: Flagging users who spread toxic content online is a significant challenge. The state $x_{k}$ represents the category of peddlers classified based on the intensity and type of content they are propagating. For example, the state could be 3-dimensional, indicating the hate-intent of the user (hateful or not), the hate speech intensity scale [90], and the particular group the hate speech is directed towards. The noisy observations are the text comments from the user that inform about the state and are from a high-dimensional observation space.
In a social network, there may be multiple LLMAs deployed to flag malicious users and decrease the propagation of hate speech. The flags by the previous LLMA are visible, but the private observations are not due to computation and privacy restrictions (so that the LLM can not be fine-trained on the text or the feature mappings). Since the observations are generated sequentially, the LLMA may use the same LLM but with a different context and for a different text observation. Therefore, a single LLMA can be viewed as a sequence of LLMAs learning from their private observations and the past actions of previous LLMAs.
@@
The first time accounts for the misclassification of a hateful user, and the second term accounts for the difference in intensity. We use the measuring hate speech dataset from [54], which contains $40,000$ annotated comments from Twitter, Reddit, and Gab. The annotations are performed by crowdsourcing and indicate if the comments contain hate speech and the intensity of the toxicity exhibited on a scale of 1 to 5, measured using a Rasch measurement scale [54].
@@ XI-B2 Product Quality Identification
In the case of an e-commerce platform, the restriction on the sharing of the private observation is motivated by the computational constraints that the LLMAs will have when performing Bayesian inference on millions of products and billions of reviews. The public belief can be efficiently computed based on the output (action) of each of the LLMA.
@@
We now detail on what we mean by quality and the definition of the states. Since we do not have access to any additional dataset other than ratings, we consider products with at least 10001000 10001000 ratings and compute the quality state ex-ante using these ratings. Therefore, the quality here is more representative of the perceived value for money [92]. We consider the following ranges $[0,3.3),[3.3,5]$ of the average review of all the ratings to assign a product quality as bad, medium, and good, respectively. To clarify, we just use all 10001000 10001000 ratings to compute the state; the observations are sampled one at a time and analyzed by the LLMA.
@@
Figure 12: Interpretability of the outputs of an LLM Sensor for Product Quality Identification. There are 44 44 different features that the LLM of the LLMA extracts. The left subplot corresponds to reviews from a good product, and the right subplot corresponds to reviews from a bad product (defined in the text). We analyze text reviews with ratings from 11 11 to 55 55 for 44 44 different attributes and plot the proportion of samples where the attribute is true. It can be seen that the features, indeed, are consistent with the overall rating provided by the reviewer for the two different kinds of products and can be used for a more fine-grained interpretable analysis.